OpenStudy (anonymous):

HELP!! Find and simplify the composite function (h o h)(x) where h(x)= (x+7)/(x-7). Thanks

OpenStudy (anonymous):

\[ (h\circ h)(x) = h(h(x)) = h\left(\frac{x+7}{x-7}\right) = h\left(\frac{\frac{x+7}{x-7}+7}{\frac{x+7}{x-7}-7}\right) \]Now you must simplify.

OpenStudy (anonymous):

yup you are right, but i am actually having a hard time simplifying this :/

OpenStudy (anonymous):

can you help me a little bit more?

OpenStudy (anonymous):

Multiply the top and bottom by \(x-7\) for starters.

OpenStudy (anonymous):

\[ \frac{\frac{x+7}{x-7}+7}{\frac{x+7}{x-7}-7}=\frac{x+7+7(x-7)}{x+7-7(x-7)} \]

OpenStudy (anonymous):

Can you do the rest from here?